what would be the answer and solution for this equation?
4x+3y=26
y=3x

Answer by Maths68(1474)

4x+3y=26..........(1)

y=3x
y-3x=0............(2)
Multiply (2) by 3 and subtract it from (1)
4x+3y=26
-3y+9x=0
-----------------------
13x=26
Divide by 13 both sides of the above equation
13x/13=26/13
x=2
Put the value of x in (2)
y=3x
y=3(2)
y=6

(x,y)=(2,6)



Check
======
Put the value of x and y in (1)
4x+3y=26..........(1)
4(2)+3(6)=26
8+18=26
26=26